OpenStudy (anonymous):

\[\sqrt{s^8}+\sqrt{25^8}+2\sqrt{s^8}+\sqrt{s^4}\]

OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form

OpenStudy (anonymous):

help please :)

OpenStudy (anonymous):

First combine like terms to get and change to exponential notation:\[3s ^{(8/2)}+s ^{(4/2)}+25^{(8/2)}\]

OpenStudy (anonymous):

okay...yea?

OpenStudy (anonymous):

Simplify to get: \[3s ^{4}+s ^{2}+390625\]

OpenStudy (anonymous):

That last number is huge, are you sure it was the square root of 25 to the 8th?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

yea 25s^8

OpenStudy (anonymous):

Okay... you need to factor after using a u substitution for s^2 and u^2=s^4 giving: \[3u ^{2}+u +390625\] Make sure they are regular square roots to start and not 4th roots too.

OpenStudy (anonymous):

I'm going too far... if this is not set to zero, we can't solve for s... the answer above is your simplest form (before the u substitution):\[3s ^{4}+s ^{2}+390625\]

OpenStudy (anonymous):

but thats not in the answer choice

OpenStudy (swissgirl):

noo she wrote it incorrect

OpenStudy (anonymous):

lol i did?

OpenStudy (anonymous):

s^4 + 25^4 + 2s^4 + s^2

OpenStudy (anonymous):

more to do

OpenStudy (swissgirl):

\( \large\sqrt{s^8} + \sqrt{25s^8} + s\sqrt{s^8} +\sqrt{s^4} \)

OpenStudy (anonymous):

yes ^^

OpenStudy (swissgirl):

if you want the answer i can give it to u :P lol \( s^4+5s^4+2s^4+s^2 \) \( 8s^4+s^2 \) That is your answer

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

lol, well that makes a difference

OpenStudy (anonymous):

there is a missing s in the second term!!

OpenStudy (anonymous):

According to the answer that Swissgirl gave, her original statement of the true question was also incorrect, but very close. One of the s' should be a two: \[\sqrt{s ^{8}}+\sqrt{25s ^{8}}+2\sqrt{s ^{8}}+\sqrt{s ^{4}}\]then go to Swissgirl's final solution.

OpenStudy (swissgirl):

ya it was an accident it shld have been a 2 I had solved this question b4 but I guess the user didnt follow my explanation so i just decided to give the answer

OpenStudy (anonymous):

ty